OpenStudy (anonymous):

Solve for y: z ___________ = x (4 + Y)

OpenStudy (anonymous):

multiply both sides by (4 + y) z = x(4 + y) distribute z = 4x + xy subtract 4x z - 4x = xy divide by x z - 4x ------ = y x